Find the approximate angle between vectors: 𝐒 = (2, 3, 5) and 𝐓⃗ = (1, 6, -4).

1
Expert's answer
2022-05-30T15:18:43-0400

"\alpha=\arccos\frac{2\cdot 1+3\cdot 6-5\cdot 4}{\sqrt{2^2+3^2+5^2}\cdot \sqrt{1^2+6^2+4^2}}=90°."